how to find distance in 2nd degree form of pair of straight line. there is any formula.

Grade:10

1 Answers

Vikas TU
6 years ago
Dear Student,
There is a formula to find the distance between a pair of parallel lines, if the general equation is given.
ax2+ 2hxy + by2+ 2fy + c = 0 represents a pair of parallel straight lines if h/a = b/h = f/g or bg2 = af2
The distance between them is given by 2√g2–ac/a(a+b) or √f2–bc/b(a+b).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)